OpenStudy (anonymous):

write an equation passing through (13,9) (30,26)

OpenStudy (anonymous):

well, the slope of (13,9) and (30,26) is 7/17. You can write it in point slope form. A.K.A y-y=m(x-x) so... \[y-9= 7/17 (x-13)\]